OpenStudy (anonymous):

Ava, Grace, Jayden, and Zoe together stacked a total of 50 brownies on a tray. Ava stacked 0.14 of the brownies, Grace stacked 4 of the 50 brownies, Jayden stacked 24% of the brownies, and Zoe stacked the remaining brownies. Who stacked the most brownies?

OpenStudy (michele_laino):

please in order to solve your problem, keep in mind that \[0.14=\frac{ 7 }{ 50 }\] furthermore to each fraction corresponds a number of brownies, namely to 7/50, corresponds \[\frac{ 7 }{ 50 }*50=7\] brownies. Please now do the same using other fraction numbers

OpenStudy (michele_laino):

hint: to 24% corresponds this fraction \[\frac{ 24 }{ 100 }=\frac{ 12 }{ 50 }\]

OpenStudy (michele_laino):

so how many brownies correspond to 12/50? please try

OpenStudy (anonymous):

12 @Michele_Laino

OpenStudy (michele_laino):

perfect, now we have the subsequent numbers of brownies: Ava =7 Grace =4 Jayden = 12 and Zoe?

OpenStudy (michele_laino):

please note that sum of Ava+Grace+Jayden+Zoe =50 so?

OpenStudy (anonymous):

17? @Michele_Laino

OpenStudy (michele_laino):

@abbyfox101 please try

OpenStudy (michele_laino):

are you sure? 7+4+12+ 17 is equals to 50?

OpenStudy (michele_laino):

@abbyfox101 please retry, take your time

OpenStudy (anonymous):

27? @Michele_Laino

OpenStudy (michele_laino):

@abbyfox101 well done! so your answer is?

OpenStudy (anonymous):

zoe @Michele_Laino

OpenStudy (michele_laino):

@abbyfox101 that's right! Congratulations!!!

OpenStudy (anonymous):

Yay! Thanks! @Michele_Laino

OpenStudy (michele_laino):

@abbyfox101 Thanks!